OpenStudy (anonymous):

The measure of the angle is 17 times greater than it supplement ?? Need help

OpenStudy (phi):

call the angle x is supplement is 180-x x is 17 times bigger than (180-x) x = 17(180-x) x= 17*180 - 17x add 17 x to both sides 18x= 17*180 divide both sides by 18 x= 17*180/18= 17*10= 170